# arXiv:2603.01498v1[cs.CV] 02 Mar 2026

### v1

@@ Abstract
1 Introduction
Remote sensing change detection, which aims to identify regions of interest that have changed between bi-temporal satellite images, plays an irreplaceable role in numerous scenarios such as urban environmental monitoring, agricultural surveillance, disaster assessment, and damage evaluation Ding et al. (2025). However, Binary Change Detection (BCD) Marchesi and Bruzzone (2009) merely answers “where changes have occurred,” offering limited actionable information that is often insufficient to guide follow-up actions. Semantic Change Detection (SCD) Ding et al. (2024a), on the other hand, requires performing full semantic segmentation on both images before comparing them, which is computationally intensive and demands exhaustive annotation of all semantic categories in both temporal phases. For example, in post-disaster damage assessment, knowing only that “changes have occurred” is far from adequate. It is also essential to clarify “what the changes signify” and to complete the assessment rapidly. Therefore, under real-world constraints, a balance must be struck between efficiency and performance. Multi-class change detection (MCD) Zhu et al. (2024) addresses this need by subdividing changes into specific semantic categories, thereby providing finer-grained information while significantly reducing annotation effort. This makes MCD a suitable solution for time-sensitive tasks such as rapid damage assessment.
Despite its practical potential, the technical implementation of MCD remains challenging. In real-world scenarios, changed areas are typically small, occupying only a fraction of the total image pixels. Furthermore, due to factors such as damage or degradation, the features distinguishing different change categories are often subtle, making accurate classification particularly difficult. Traditional approaches, including deep learning methods based on siamese networks, commonly suffer from performance degradation due to insufficient information. Additionally, while powerful self-supervised foundation models such as DINOv3 Siméoni et al. (2025) are capable of extracting rich, hierarchical features, they have not yet been fully explored or adapted for this specific task. Direct application often fails to explicitly model and integrate the complementary features needed to distinguish fine-grained inter-class differences within changed regions, such as variations in structural damage severity.
@@ 2 Related Work
2.1 Remote Sensing Change Detection
@@
BCD aims to produce a binary mask that labels each pixel as “changed” or “unchanged.” Early approaches relied on traditional machine learning and image processing techniques Marchesi and Bruzzone (2009). With the deep learning revolution, siamese architectures based on convolutional neural networks have become mainstream. For example, the FC-Siam network Daudt et al. (2018); Zhan et al. (2017) employs weight-sharing to extract features from bi-temporal images. To enhance the feature extraction capability of convolutional networks, Fang et al. proposed SNUNet Fang et al. (2021a), which adopts dense skip connections in a U-shaped architecture. Furthermore, the inherently local receptive field of CNNs limits their ability to model long-range dependencies and capture global contextual information. To address this issue, researchers have introduced the attention mechanism and Transformer architecture Vaswani et al. (2017); Dosovitskiy et al. (2021). He et al. proposed a channel bias split attention (CBSA) to recover the information in the change region by adding a simple channel mapping He et al. (2024). The Transformer architecture establishes dynamic interactions among all elements within an image by utilizing a self-attention mechanism. This enables the model to explicitly model global dependencies and long-range contextual relationships, which is crucial for distinguishing genuine changes from spurious variations across different regions. In Chen et al. (2021), a hybrid architecture combining Transformer and CNN, termed the BIT model, was proposed. This model employs a Transformer structure to process features obtained via convolution. In contrast, ChangeFormer Bandara and Patel (2022) adopts a pure Transformer framework. Zhang et al. introduced SwinSUNet Zhang et al. (2022), which incorporates the Swin Transformer Liu et al. (2021b) structure into the BCD task. These studies demonstrate the critical role of advanced network architecture design in feature extraction.
@@
Both MCD Zhu et al. (2024) and SCD are inherently segmentation tasks, requiring higher-level semantic understanding compared to BCD. Due to the demand for fine-grained semantic annotations, most publicly available datasets are constructed under the BCD paradigm, resulting in relatively limited research and data resources dedicated to MCD and SCD. MCD aims to classify changed pixels into specific semantic categories. SCD, by nature, is a bi-temporal semantic segmentation task that requires generating precise semantic maps for each temporal phase. Owing to their fundamental similarities, the distinction between MCD and SCD is often not strictly emphasized in the literature Cui and Jiang (2023). Early research in SCD, such as Mou et al. (2018), introduced an end-to-end recurrent convolutional neural network (ReCNN) that integrates the spectral-spatial feature extraction capabilities of CNNs with the temporal modeling strengths of RNNs. Long et al. proposed a multi-task learning framework that decomposes the detection of change regions into three sub-tasks: bi-temporal semantic segmentation, change detection, and boundary detection Long et al. (2025). Ding et al. introduced SCanNet, which designs a semantic learning loss based on temporal consistency constraints Ding et al. (2024a). Similarly, in Tang et al. (2024), BCD is also incorporated as a subtask, and a contrastive learning module is designed to interact information between semantic detection and change detection.
2.2 Visual Foundation Models
@@
Visual Foundation Models (VFMs) acquire powerful general visual representation capabilities through pre-training on large-scale, multi-source datasets. This opens new perspectives for visual change detection. CLIP Radford et al. (2021) learns from image-text pairs to align visual and textual representations, enabling open-vocabulary reasoning. Segment Anything Model (SAM) Kirillov et al. (2023) demonstrates exceptional zero-shot generalization capability on segmentation tasks. In contrast, the DINO series Siméoni et al. (2025) relies solely on self-supervised learning to extract general-purpose visual features from massive datasets, making it well-suited as a feature encoder.
@@
The success of VFMs has naturally attracted significant interest from researchers, and how to adapt and inject domain-specific knowledge into these foundation models has become a current focus. A common paradigm is to adopt SAM as the backbone, design it in a siamese architecture, and then train it on the target domain Qin et al. (2025). Furthermore, fine-tuning these VFMs has led to further performance improvements in change detection. For instance, Gao et al. (2025) designed CNN adapters for this purpose. Meanwhile, Zhang et al. (2024) introduced a parameter-efficient fine-tuning approach using LoRA-like adapters. To mitigate the substantial parameter overhead of VFMs, Ding et al. (2024b) alternatively employed lightweight pre-trained variants such as FastSAM Zhao et al. (2023) and MobileSAM Zhang et al. (2023).
@@
Despite the powerful representational capabilities of models like SAM, their robustness on remote sensing images remains limited, often struggling to identify RS-specific targets comprehensively. This significantly hinders the application of VFMs in SCD and MCD tasks. Mei et al. (2024) first introduced SAM into the SCD task, meticulously designing a dual encoder-decoder architecture. This approach cross-fuses features from MobileSAM and a CNN encoder, and employs a multi-task scheme in the decoder to progressively decode features at different levels. Jiang et al. (2025) proposed a large-scale remote sensing dataset for fine-tuning and simultaneously utilized both CLIP and SAM to encode bi-temporal features. Zheng et al. (2025) was the first to incorporate DINO as a feature extractor for MCD, also adopting a multi-scale feature fusion architecture to enhance detection performance. In summary, effectively leveraging the potential of VFMs for high-level semantic tasks like SCD and MCD necessitates deliberate design of data and models.
3 Method
@@ 3.4 Enchaned Decoder
3.5 Loss Function
In multi-class change detection for remote sensing imagery, the regions of interest (i.e., changed areas) typically occupy only a small fraction of the image. Moreover, severe class imbalance exists among different change categories. To mitigate these training challenges, various weighted loss functions have been proposed. In this work, we adopt a composite loss combining Focal Loss Lin et al. (2017), Dice Loss Milletari et al. (2016), and Lovász Loss Berman et al. (2018) to jointly address the class-imbalance issue. Focal Loss down-weights well-classified examples and focuses on hard samples; Dice Loss directly optimizes the overlap between predictions and ground truth; and Lovász Loss, as a differentiable surrogate for the Jaccard index, helps refine segmentation boundaries. The combination of these losses ensures stable gradient propagation and encourages the model to achieve balanced performance across all change categories. Then, we have:
@@ 3.5 Loss Function
4 Experiments
4.1 Datasets
SECOND-CD Yang et al. (2021) is a large-scale, high-quality dataset specifically constructed for semantic change detection. The imagery covers three major Chinese cities: Hangzhou, Chengdu, and Shanghai, comprising a total of 4,662 paired high-resolution aerial/satellite images with a size of 512×512 pixels (spatial resolution 0.3–5 m). The dataset is split into 2,968 pairs for training and 1,694 pairs for testing. It provides bi-temporal semantic labels and change semantic labels, encompassing categories such as non-vegetated ground surface, tree, low vegetation, water, buildings, and playgrounds, along with their corresponding change information.
Gaza-change Zheng et al. (2025) is a dataset for multi-class change detection in the Gaza Strip. It comprises 922 precisely co-registered image pairs, each of 512×512 pixels, captured by the Beijing-2 satellite (3.2 m Ground Sampling Distance) between 2023 and 2024. The imagery covers nine major urban areas, including Khan Yunis and Rafah. Following the annotation paradigm of MCD tasks, where only changes are annotated. The dataset labels the target pixel into one of six semantic change categories: building damage, new building, new camp, farmland damage, greenhouse damage, and new greenhouse. The dataset is randomly split into 554 pairs for training, 184 for validation, and 184 for testing.
4.2 Implementation Details
All experiments in this paper were conducted on a single GPU (Nvidia RTX 3090 Ti) using the PyTorch framework. The models were trained for 150 epochs, with validation performed at the end of each epoch. We employed the AdamW optimizer with an initial learning rate of 0.0001 and a weight‑decay coefficient of 0.05. The coefficients $\alpha,\beta,\gamma$ in the loss function are set to 0.4, 0.3, and 0.3, respectively. Furthermore, to adapt the comparative methods to the MCD task, the output categories of these methods were correspondingly modified. It should be noted that the SECOND dataset provides pixel‑wise labels only for changed regions and does not supply separate bi‑temporal semantic labels. We evaluated our method by comparing it with a range of state-of-the-art techniques that represent diverse methodological lines. The selected baselines include: BIT Chen et al. (2021), SNUNet Fang et al. (2021b), LGPNet Liu et al. (2021a), ChangeFormer Bandara and Patel (2022), USSFCNet Lei et al. (2023), DDLNet Ma et al. (2024), Rsmamba Liu et al. (2024), CDxLstm Wu et al. (2025), AnyChange Zheng et al. (2024), and MC-DiSNet Zheng et al. (2025).
4.3 Evaluation Metrics
To comprehensively evaluate the performance of MCD tasks, we adopt the four metrics proposed in Ding et al. (2022) (identical to those used for SCD): Overall Accuracy (OA), mean Intersection over Union (mIoU), Separated Kappa (SeK), and the SCD variant of the F1-score (denoted as $F_{SCD}$). OA is defined as follows:
@@ 4.4 Main Results
As shown in Figure 2 and Figure 3, we visually compare the detection results of different methods on the Gaza-change dataset. Tri-path DINO produces much clearer detection boundaries, demonstrating that the proposed complementary learning architecture enables the model to better understand the detailed features of changes. Simultaneously, by examining the first row of results in both figures, it can be observed that the proposed method has a very low false detection rate, which proves Tri-path DINO’s high detection performance for small target areas. For larger targets, Tri-path achieves the most complete detection. This ability to adapt to targets of varying sizes strongly indicates the potential of the complementary learning framework in remote sensing tasks.
As reported in Table 1, our proposed Tri-path DINO achieves the best performance across all four evaluation metrics. Specifically, our model attains an OA of 96.45% and an mIoU of 71.35%, representing improvements of 0.12% and 0.71%, respectively, over the previous best method, MC-DiSNet. This validates the superiority of our model in terms of overall classification accuracy and segmentation consistency across various change categories. More importantly, on the Sek and $F_{scd}$ metrics, which better reflect a model’s ability to distinguish between imbalanced change classes, our method achieves significant leads of 13.02 and 62.58, substantially outperforming all other compared methods.
@@
Table 1: Comprehensive comparison with state-of-the-art methods on the Gaza-change dataset.
@@
Table 2: Comprehensive comparison with state-of-the-art methods on the SECOND dataset.
We compared various methods on the classic SECOND dataset. Table 2 shows that Tri-path DINO achieves highly competitive performance on this standard benchmark. Notably, our method attains the highest scores in three key metrics: mIoU (71.62%), Sek (20.61), and $F_{scd}$ (55.44). Compared to the previous state-of-the-art method, MC-DiSNet, Tri-path DINO outperforms it across all four evaluated metrics, demonstrating the effectiveness of the complementary feature learning strategy. In contrast, the AnyChange method, based on a pre-trained foundational model, performs significantly below expectations due to its difficulty in handling shifts in input distribution, highlighting the necessity of our fine-tuning strategy.
4.5 Ablation Study

### v2

# Tri-path DINO: Feature Complementary Learning for Remote Sensing Multi-Class Change Detection

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
Remote sensing change detection, which aims to identify regions of interest that have changed between bi-temporal satellite images, plays an irreplaceable role in numerous scenarios such as urban environmental monitoring, agricultural surveillance, disaster assessment, and damage evaluation [7]. However, Binary Change Detection (BCD) [24] merely answers “where changes have occurred,” offering limited actionable information that is often insufficient to guide follow-up actions. Semantic Change Detection (SCD) [8], on the other hand, requires performing full semantic segmentation on both images before comparing them, which is computationally intensive and demands exhaustive annotation of all semantic categories in both temporal phases. For example, in post-disaster damage assessment, knowing only that “changes have occurred” is far from adequate. It is also essential to clarify “what the changes signify” and to complete the assessment rapidly. Therefore, under real-world constraints, a balance must be struck between efficiency and performance. Multi-class change detection (MCD) [42] addresses this need by subdividing changes into specific semantic categories, thereby providing finer-grained information while significantly reducing annotation effort. This makes MCD a suitable solution for time-sensitive tasks such as rapid damage assessment.
Despite its practical potential, the technical implementation of MCD remains challenging. In real-world scenarios, changed areas are typically small, occupying only a fraction of the total image pixels. Furthermore, due to factors such as damage or degradation, the features distinguishing different change categories are often subtle, making accurate classification particularly difficult. Traditional approaches, including deep learning methods based on siamese networks, commonly suffer from performance degradation due to insufficient information. Additionally, while powerful self-supervised foundation models such as DINOv3 [30] are capable of extracting rich, hierarchical features, they have not yet been fully explored or adapted for this specific task. Direct application often fails to explicitly model and integrate the complementary features needed to distinguish fine-grained inter-class differences within changed regions, such as variations in structural damage severity.
@@ II Related Work
II-A Remote Sensing Change Detection
@@
BCD aims to produce a binary mask that labels each pixel as “changed” or “unchanged.” Early approaches relied on traditional machine learning and image processing techniques [24]. With the deep learning revolution, siamese architectures based on convolutional neural networks have become mainstream. For example, the FC-Siam network [5, 35] employs weight-sharing to extract features from bi-temporal images. To enhance the feature extraction capability of convolutional networks, Fang et al. proposed SNUNet [11], which adopts dense skip connections in a U-shaped architecture. Furthermore, the inherently local receptive field of CNNs limits their ability to model long-range dependencies and capture global contextual information. To address this issue, researchers have introduced the attention mechanism and Transformer architecture [32, 10]. He et al. proposed a channel bias split attention (CBSA) to recover the information in the change region by adding a simple channel mapping [14]. The Transformer architecture establishes dynamic interactions among all elements within an image by utilizing a self-attention mechanism. This enables the model to explicitly model global dependencies and long-range contextual relationships, which is crucial for distinguishing genuine changes from spurious variations across different regions. In [3], a hybrid architecture combining Transformer and CNN, termed the BIT model, was proposed. This model employs a Transformer structure to process features obtained via convolution. In contrast, ChangeFormer [1] adopts a pure Transformer framework. Zhang et al. introduced SwinSUNet [37], which incorporates the Swin Transformer [21] structure into the BCD task. These studies demonstrate the critical role of advanced network architecture design in feature extraction.
@@
Both MCD [42] and SCD are inherently segmentation tasks, requiring higher-level semantic understanding compared to BCD. Due to the demand for fine-grained semantic annotations, most publicly available datasets are constructed under the BCD paradigm, resulting in relatively limited research and data resources dedicated to MCD and SCD. MCD aims to classify changed pixels into specific semantic categories. SCD, by nature, is a bi-temporal semantic segmentation task that requires generating precise semantic maps for each temporal phase. Owing to their fundamental similarities, the distinction between MCD and SCD is often not strictly emphasized in the literature [4]. Early research in SCD, such as [27], introduced an end-to-end recurrent convolutional neural network (ReCNN) that integrates the spectral-spatial feature extraction capabilities of CNNs with the temporal modeling strengths of RNNs. Long et al. proposed a multi-task learning framework that decomposes the detection of change regions into three sub-tasks: bi-temporal semantic segmentation, change detection, and boundary detection [22]. Ding et al. introduced SCanNet, which designs a semantic learning loss based on temporal consistency constraints [8]. Similarly, in [31], BCD is also incorporated as a subtask, and a contrastive learning module is designed to interact information between semantic detection and change detection.
II-B Visual Foundation Models
@@
Visual Foundation Models (VFMs) acquire powerful general visual representation capabilities through pre-training on large-scale, multi-source datasets. This opens new perspectives for visual change detection. CLIP [29] learns from image-text pairs to align visual and textual representations, enabling open-vocabulary reasoning. Segment Anything Model (SAM) [16] demonstrates exceptional zero-shot generalization capability on segmentation tasks. In contrast, the DINO series [30] relies solely on self-supervised learning to extract general-purpose visual features from massive datasets, making it well-suited as a feature encoder.
@@
The success of VFMs has naturally attracted significant interest from researchers, and how to adapt and inject domain-specific knowledge into these foundation models has become a current focus. A common paradigm is to adopt SAM as the backbone, design it in a siamese architecture, and then train it on the target domain [28]. Furthermore, fine-tuning these VFMs has led to further performance improvements in change detection. For instance, [13] designed CNN adapters for this purpose. Meanwhile, [38] introduced a parameter-efficient fine-tuning approach using LoRA-like adapters. To mitigate the substantial parameter overhead of VFMs, [9] alternatively employed lightweight pre-trained variants such as FastSAM [39] and MobileSAM [36].
@@
Despite the powerful representational capabilities of models like SAM, their robustness on remote sensing images remains limited, often struggling to identify RS-specific targets comprehensively. This significantly hinders the application of VFMs in SCD and MCD tasks. [25] first introduced SAM into the SCD task, meticulously designing a dual encoder-decoder architecture. This approach cross-fuses features from MobileSAM and a CNN encoder, and employs a multi-task scheme in the decoder to progressively decode features at different levels. [15] proposed a large-scale remote sensing dataset for fine-tuning and simultaneously utilized both CLIP and SAM to encode bi-temporal features. [40] was the first to incorporate DINO as a feature extractor for MCD, also adopting a multi-scale feature fusion architecture to enhance detection performance. In summary, effectively leveraging the potential of VFMs for high-level semantic tasks like SCD and MCD necessitates deliberate design of data and models.
III Method
@@ III-D Enchaned Decoder
III-E Loss Function
In multi-class change detection for remote sensing imagery, the regions of interest (i.e., changed areas) typically occupy only a small fraction of the image. Moreover, severe class imbalance exists among different change categories. To mitigate these training challenges, various weighted loss functions have been proposed. In this work, we adopt a composite loss combining Focal Loss [18], Dice Loss [26], and Lovász Loss [2] to jointly address the class-imbalance issue. Focal Loss down-weights well-classified examples and focuses on hard samples; Dice Loss directly optimizes the overlap between predictions and ground truth; and Lovász Loss, as a differentiable surrogate for the Jaccard index, helps refine segmentation boundaries. The combination of these losses ensures stable gradient propagation and encourages the model to achieve balanced performance across all change categories. Then, we have:
@@ III-E Loss Function
IV Experiments
IV-A Datasets
SECOND-CD [34] is a large-scale, high-quality dataset specifically constructed for semantic change detection. The imagery covers three major Chinese cities: Hangzhou, Chengdu, and Shanghai, comprising a total of 4,662 paired high-resolution aerial/satellite images with a size of 512×512 pixels (spatial resolution 0.3–5 m). The dataset is split into 2,968 pairs for training and 1,694 pairs for testing. It provides bi-temporal semantic labels and change semantic labels, encompassing categories such as non-vegetated ground surface, tree, low vegetation, water, buildings, and playgrounds, along with their corresponding change information.
Gaza-change [40] is a dataset for multi-class change detection in the Gaza Strip. It comprises 922 precisely co-registered image pairs, each of 512×512 pixels, captured by the Beijing-2 satellite (3.2 m Ground Sampling Distance) between 2023 and 2024. The imagery covers nine major urban areas, including Khan Yunis and Rafah. Following the annotation paradigm of MCD tasks, where only changes are annotated. The dataset labels the target pixel into one of six semantic change categories: building damage, new building, new camp, farmland damage, greenhouse damage, and new greenhouse. The dataset is randomly split into 554 pairs for training, 184 for validation, and 184 for testing.
IV-B Implementation Details
All experiments in this paper were conducted on a single GPU (Nvidia RTX 3090 Ti) using the PyTorch framework. The models were trained for 150 epochs, with validation performed at the end of each epoch. We employed the AdamW optimizer with an initial learning rate of 0.0001 and a weight‑decay coefficient of 0.05. The coefficients $\alpha,\beta,\gamma$ in the loss function are set to 0.4, 0.3, and 0.3, respectively. Furthermore, to adapt the comparative methods to the MCD task, the output categories of these methods were correspondingly modified. It should be noted that the SECOND dataset provides pixel‑wise labels only for changed regions and does not supply separate bi‑temporal semantic labels. We evaluated our method by comparing it with a range of state-of-the-art techniques that represent diverse methodological lines. The selected baselines include: BIT [3], SNUNet [12], LGPNet [20], ChangeFormer [1], USSFCNet [17], DDLNet [23], Rsmamba [19], CDxLstm [33], AnyChange [41], and MC-DiSNet [40].
IV-C Evaluation Metrics
To comprehensively evaluate the performance of MCD tasks, we adopt the four metrics proposed in [6] (identical to those used for SCD): Overall Accuracy (OA), mean Intersection over Union (mIoU), Separated Kappa (SeK), and the SCD variant of the F1-score (denoted as $F_{SCD}$). OA is defined as follows:
@@ IV-D Main Results
Figure 4: Example results on SECOND dataset.
Figure 5: Qualitative comparison on the SECOND dataset, with color coding: true positives (white), true negatives (black), false positives (red), and false negatives (green).
As shown in Figure 2 and Figure 3, we present a visual comparison of the detection results of different methods on the Gaza-change dataset. Tri-path DINO produces clearer detection boundaries, indicating that the proposed complementary learning architecture enables the model to better understand the detailed features of changes. Meanwhile, by observing the results in the first row of Figure 2 and the second row of Figure 3, it can be seen that the proposed method improves the recognition rate of small targets and avoids the false detections encountered by other methods, demonstrating the high detection performance of Tri-path DINO for fine-grained features. For larger targets, Tri-path achieves the most complete detection. This ability to adapt to targets of varying sizes strongly demonstrates the potential of the complementary learning framework in remote sensing tasks.
Compared with the facility damage, the SECOND dataset contains more detailed variations. Figure 4 and Figure 5 present the detection results of various state-of-the-art methods on this dataset. It can be clearly observed that the proposed Tri-path DINO captures finer details. In particular, the gaps between buildings in the third row are identified, a capability not achieved by other methods. Additionally, the subtle changes in the central area of the fourth row are also recognized, reflecting the sensitivity of complementary learning to changes. In summary, the complementary learning framework significantly enhances the discriminative ability of the pre-trained model for fine-grained features.
As reported in Table I, our proposed Tri-path DINO achieves the best performance across all four evaluation metrics. Specifically, our model attains an OA of 96.45% and an mIoU of 71.35%, representing improvements of 0.12% and 0.71%, respectively, over the previous best method, MC-DiSNet. This validates the superiority of our model in terms of overall classification accuracy and segmentation consistency across various change categories. More importantly, on the Sek and $F_{scd}$ metrics, which better reflect a model’s ability to distinguish between imbalanced change classes, our method achieves significant leads of 13.02 and 62.58, substantially outperforming all other compared methods.
@@
TABLE I: Comprehensive comparison with state-of-the-art methods on the Gaza-change dataset.
@@
TABLE II: Comprehensive comparison with state-of-the-art methods on the SECOND dataset.
We compared various methods on the classic SECOND dataset. Table II shows that Tri-path DINO achieves highly competitive performance on this standard benchmark. Notably, our method attains the highest scores in three key metrics: mIoU (71.62%), Sek (20.61), and $F_{scd}$ (55.44). Compared to the previous state-of-the-art method, MC-DiSNet, Tri-path DINO outperforms it across all four evaluated metrics, demonstrating the effectiveness of the complementary feature learning strategy. In contrast, the AnyChange method, based on a pre-trained foundational model, performs significantly below expectations due to its difficulty in handling shifts in input distribution, highlighting the necessity of our fine-tuning strategy.
IV-E Ablation Study